OpenStudy (anonymous):

Solve, assuming a linear equation fits the situation. The value of a graphing calculator is $225. After 2 years, the value of this calculator is $160. Find the value of the calculator after 5 years. ( Hint, Let the time now be 0, the time in two years be 2, and the time in five years be 5.)

OpenStudy (binary3i):

when t=0 cost c=225$ when t=2 c= 160$ c= (32.5)t + 95 now put t=5 you'll get your answer

OpenStudy (anonymous):

So the answer is 257.5?

OpenStudy (anonymous):

So would that be it then?

OpenStudy (dumbcow):

no each year the cost is decreasing by 32.5 c(t) = 225 -32.5t

OpenStudy (anonymous):

So the answer is 192.5?

OpenStudy (binary3i):

oh yeah made a mistake sorry.

OpenStudy (dumbcow):

plug in t=5 --> c = 225 - (32.5)(5)

OpenStudy (anonymous):

So its 62.5?

OpenStudy (dumbcow):

yes